Select the best possible first step to solving the system by first eliminating the x variable
3x + 5y = 1
2x + 4y = –4

Answer by Alan3354(69443)

Select the best possible first step to solving the system by first eliminating the x variable
3x + 5y = 1
2x + 4y = –4
-----------
"best possible" can be a matter of opinion.
I would multiply the 2nd equation by 1.5 --> 3x + 6y = -6